Star Videos, Inc., produces short musical videos for sale to retail outlets. The company’s balance sheet accounts as of January 1 are given below.

Star Videos, Inc.
Balance Sheet
January 1
Assets
Cash $ 90,800
Accounts receivable 104,200
Inventories:
Raw materials (film, costumes) $ 26,400
Videos in process 51,000
Finished videos awaiting sale 87,800 165,200
Prepaid insurance 8,550
Studio and equipment (net) 568,000
Total assets $ 936,750
Liabilities and Stockholders’ Equity
Accounts payable $ 244,000
Retained earnings 692,750
Total liabilities and stockholders’ equity $ 936,750

Because the videos differ in length and in complexity of production, the company uses a job-order costing system to determine the cost of each video produced. Studio (manufacturing) overhead is charged to videos on the basis of camera-hours of activity. The company’s predetermined overhead rate for the year ($40 per camera-hour) is based on a cost formula that estimated $280,000 in manufacturing overhead for an estimated allocation base of 7,000 camera-hours. Any underapplied or overapplied overhead is closed to cost of goods sold. The following transactions were recorded for the year:

  1. Film, costumes, and similar raw materials purchased on account, $214,000.
  2. Film, costumes, and other raw materials issued to production, $230,000 (85% of this material was considered direct to the videos in production, and the other 15% was considered indirect).
  3. Utility costs incurred (on account) in the production studio, $91,800.
  4. Depreciation recorded on the studio, cameras, and other equipment, $96,400. Three-fourths of this depreciation related to actual production of the videos, and the remainder related to equipment used in marketing and administration.
  5. Advertising expense incurred (on account), $149,000.
  6. Salaries and wages paid in cash as follows:
Direct labor (actors and directors) $ 90,400
Indirect labor (carpenters to build sets, costume designers, and so forth) $ 83,500
Administrative salaries $ 113,200
  1. Prepaid insurance expired during the year, $7,800 (70% related to production of videos, and 30% related to marketing and administrative activities).
  2. Miscellaneous marketing and administrative expenses incurred (on account), $9,900.
  3. Studio (manufacturing) overhead was applied to videos in production. The company recorded 7,250 camera-hours of activity during the year.
  4. Videos that cost $552,000 to produce according to their job cost sheets were transferred to the finished videos warehouse to await sale and shipment.
  5. Sales for the year totaled $1,110,000 and were all on account.
  6. The total cost to produce the videos that were sold according to their job cost sheets was $596,090.
  7. Collections from customers during the year totaled $1,060,000.
  8. Payments to suppliers on account during the year, $610,000.
  9. Underapplied or overapplied overhead $__?__.

Required:

1. Prepare a transaction analysis that records all of the above transactions.

2. Prepare a schedule of cost of goods manufactured for the year.

3. Prepare a schedule of cost of goods sold for the year.

4. Prepare an income statement for the year.
Answer #1

1.

Star Videos, Inc.
Transaction Analysis
For the Year Ended December 31
Cash Accounts Receivable Raw Materials Videos in Process Finished Videos Manufacturing Overhead Prepaid Insurance Studio & Equipment (net) = Accounts Payable Retained Earnings
Beginning balance @ 1/1 90800 104200 26400 51000 87800 0 8550 568000 = 244000 692750
a) Raw material purchases 214000 = 214000
b) Raw materials used -230000 195500 34500 =
c) Utility costs 91800 = 91800
d) Depreciation charges 72300 -96400 = -24100
e) Advertising = 149000 -149000
f) Salaries & wages -287100 90400 83500 = -113200
g) Prepaid insurance 5460 -7800 = -2340
h) Miscellaneous marketing = 9900 -9900
i) Applied overhead 290000 -290000 =
j) Transfer completed videos to finished goods -552000 552000 =
k) Sales 1110000 = 1110000
l) Transfer finished goods to cost of goods sold -596090 = -596090
m) Cash collections from customers 1060000 -1060000 =
n) Payment to suppliers -610000 = -610000
o) Overapplied overhead 2440 = 2440
Ending balances @12/31 253700 154200 10400 74900 43710 0 750 471600 = 98700 910560

2.

Star Videos, Inc.
Schedule of Cost of Goods Manufactured
For the Year Ended December 31
Beginning raw materials inventory $ 26400
Add: Raw material purchases 214000
Raw materials available for use 240400
Less: Ending raw materials inventory 10400
Raw materials used 230000
Less: Indirect materials 34500 195500
Direct labor 90400
Manufacturing overhead 290000
Total manufacturing costs 575900
Add: Beginning videos in process inventory 51000
Total videos in process inventory 626900
Less: Ending videos in process inventory 74900
Cost of goods manufactured 552000

3.

Star Videos, Inc.
Schedule of Cost of Goods Sold
For the Year Ended December 31
Cost of goods manufactured 552000
Add: Beginning finished goods inventory 87800
Cost of goods available for sale 639800
Less: Ending finished goods inventory 43710
Cost of goods sold 596090
Less: Overapplied overhead 2440
Adjusted cost of goods sold $ 593650

4.

Star Videos, Inc.
Income Statement
For the Year Ended December 31
Sales revenue 1110000
Cost of goods sold 593650
Gross profit 516350
Operating expenses 298540
Net income $ 217810